Problem 6-58 (Algo) Predetermined Rates and Product Profitability: Two Stage Cost Allocation (LO 6-3, 4, 5)
Answer #1

rA) engineer: 140,000/28,000= 5

user related: 230,000/ 28,750= 8

revenue related: 171,250/ 3,425,000= 0.05 or 5%


rB)
Answer #2

rB)                                         toot                                                tix

revenue                           +1,650,000                                   +1,775,000

engineer cost                  -320,000                                        -401,250

engineer admin cost         -80,000                                        -60,000

user related admin cost    -88,000                                          -142,000

revenue admin cost         -82,500                                           -88,750

profit(answer)                 1,079,500                                         1,083,000


engineer 16,000* 5= 80,000                                           5* 12,000= 60,000

user: 8* 11,000= 88,000                                                 8* 17,750= 142,000

revenue: 5%* 1,650,000= 82,500                                 5%* 1,775,000= 88,750